The parking garage at a train station offers two monthly payment plans for customers. Both plans consist of a fixed monthly cost plus an additional cost per day. Plan A charges a fixed monthly fee of R55

 plus R21

 per day. The total cost for parking under plan B for a certain number of days is given in the table below:

Number of

days parked     Cost     

(R)3

106

10

260

18

436

20

480

 

If C

 represents the total monthly parking cost of a client and n

 represents the number of times he parks his vehicle, which ONE of the following systems of equations best represents this situation?


1
Expert's answer
2021-07-16T15:34:28-0400

Plan A


"C_A=55+21n"

Plan B


"C_B(3)=b+m(3)=106"

"C_B(10)=b+m(10)=260"

"10m-3m=260-106"

"7m=154"

"m=22"

"b+22(3)=106"

"b=40"

"C_B=40+22n"

"55+21n\geq40+22n"

"n\leq15"

If "n\leq15" days, Plan B is better.


If "n\geq15" days, Plan A is better.